### v1

@@ 1 INTRODUCTION
For the V2X-Real case, SECOND is trained solely on the synthetic dataset and evaluated on the real test set of V2X-Real-IC. The results are compared against benchmark models reported in the original V2X-Real paper. The findings show that models trained on synthetic data perform on par with, and in some cases outperform, models trained on real data—setting a new standard for the utility of synthetic datasets in roadside lidar perception tasks.
@@ 2 Related Work
Real-world Roadside Lidar Datasets: Recent research has increasingly focused on using lidar technology for sensing road users from an infrastructure perspective, leading to the development of several roadside lidar datasets. In 2021, the BAAI-VANJEE dataset [7] was introduced, capturing highway and intersection scenarios across various times of day and weather conditions. In 2022, the IPS300+ dataset [4] was released, notable for featuring over 300 road users per frame, primarily due to the high volume of pedestrians and cyclists at a major intersection in Beijing. That same year, the DAIR-V2X dataset [6] became the first large-scale, multi-view dataset designed for vehicle-infrastructure cooperative driving research. Also in 2022, LUMPI [1] was introduced as the largest dataset by frame count, collected over multiple days and weather conditions. Rope3D [8], while primarily focused on monocular 3D object detection using cameras, incorporated lidar data to improve annotation accuracy. In 2023, the TUMTraf dataset [3]—a successor to the A9 dataset [19]—was released, with a focus on non-pedestrian traffic. The V2X-Real dataset [2], introduced in 2024, was the first large-scale dataset to integrate multiple vehicles and smart infrastructure for V2X cooperative perception. That same year, Int2Sec [20] targeted urban intersections using a digital twin perspective, offering 10 scenes annotated via dual roadside lidars. HoloVIC [21] further expanded the space by providing over 100K synchronized frames with 3D object annotations across multiple holographic intersections. Most recently, in 2025, the R-LiViT dataset [5] was introduced, combining lidar, RGB, and thermal imaging from a roadside viewpoint, with a focus on pedestrians and other vulnerable road users.
@@
Synthetic Roadside Lidar Datasets: The recent surge in roadside lidar datasets reflects a growing interest in infrastructure-centric perception. However, creating large-scale, annotated lidar datasets remains an inherently expensive and labor-intensive process. This limitation restricts the scalability needed for training safety-critical models [22]. Feng et al. [23] demonstrated that introducing simulation methodologies can accelerate research by a factor of 103 to 105. In response, several synthetic lidar datasets have been proposed to help overcome these challenges.
In 2022, SynLidar [24] used a custom simulator to synthesize a large synthetic lidar dataset generated using multiple simulated scenes. In the same year, V2X-Sim and V2X-Set [25] datasets simulated comprehensive multi-agent scenarios to generate synchronized camera and lidar data that included roadside units (RSUs) into the mix for cooperative perception tasks. DOLPHINS [26] dataset included six autonomous driving scenarios involving temporarily synchronized sensor data from an ego vehicle. In 2023, the DeepAccident dataset [27] was released, providing 57K frames that include safety-critical scenarios for evaluating accident prediction models. In 2024, SynthmanticLiDAR [28] introduced a CARLA-based simulator tailored for lidar semantic segmentation, producing annotated point clouds aligned with real-world class distributions. TUMTraf Synthetic (TUMTraf-Syn) Dataset [29] included 24,000 images with depth maps, and 2D and 3D annotations in 10 object categories focusing sim2real monocular 3D object detection. Most recently, the SCOPE dataset [30], also introduced in 2024, emphasized diversity—featuring realistic sensor models, physically accurate weather conditions, a catalog of over 40 scenarios, up to 24 collaborative agents, and passive traffic.
@@
Sim-to-Real Gap Mitigation Approaches: Synthetic datasets are often created using simulators that model 3D scenes under user-defined conditions. However, the outputs frequently diverge from real data due to unmodeled environmental and sensor-specific effects, resulting in a noticeable sim-to-real domain gap. Addressing this gap has been a central focus in many recent studies.
@@
3 UrbanTwin Dataset Generation and Description
We present three synthetic roadside lidar datasets: UrbanTwin-LUMPI [36], UrbanTwin-V2X-Real [37], and UrbanTwin-TUMTraf-I [38]. For brevity, these datasets will be referred to using the UT- prefix throughout this article. Each synthetic dataset is designed to replicate the core characteristics of its corresponding real-world counterpart—LUMPI [1], V2X-Real [2], and TUMTraf-I [3].
3.1 Generation Approach
@@ 3.2 Dataset Description
The synthetic datasets contain 6 classes: Car, Van, Bicycle, Motorcycle, Bus, and Truck. Due to the challenging conformation of pedestrian models in CARLA to real data, all pedestrian classes are omitted for now, and are planned to be added in a future update.
@@ 4.2 Utility in Perception Models
The high-fidelity synthetic replicas are designed not only to provide realistic simulations of real-world roadside lidar datasets but also to serve as practical resources for training and testing 3D perception models. The dataset labels are provided in a modified OpenPCDet [39] format, which includes object position (x, y, z), size (dx, dy, dz), and rotation (around the upward z-axis), with an additional column to store the object_id for use in tracking algorithm research. Segmentation labels follow the SemanticKITTI .label format [9], supporting scene understanding tasks.
@@
Case Study - 3D Object Detection: To demonstrate real-world applicability, we evaluate a common perception task—3D object detection using the synthetic datasets. The objective of this case study is not to benchmark model performance against state-of-the-art methods or to analyze detection architectures in detail. Rather, it serves as a validation exercise to assess the quality of point clouds, labels, and the overall training utility of the UT-datasets.
To this end, we train two off-the-shelf deep 3D object detectors for the car category, SEED [16] and SECOND [17], on the synthetic UT-LUMPI and UT-V2X-Real datasets, respectively, and evaluate their performance on the corresponding real test sets from LUMPI and V2X-Real.
@@
V2X-Real vs. UT-V2X-Real: For V2X-Real, we evaluate the generalization capacity of the synthetic dataset by training the SECOND model exclusively on 8,000 synthetic samples from UT-V2X and compare the results against published benchmarks from the original paper.
@@ 6 Conclusion & Future Work
This paper introduced three high-fidelity synthetic roadside lidar datasets, UT-LUMPI, UT-V2X-Real, and UT-TUMTraf-I, constructed using a digital-twin-based pipeline to replicate the physical layout and sensor properties of real-world deployments. Unlike generic synthetic datasets, our approach achieves close structural, statistical, and sensory alignment with real data. Through distributional analysis and object detection case studies, we demonstrated that models trained solely on synthetic data generalize well to real-world benchmarks, in some cases outperforming models trained on human-annotated datasets. These results highlight the utility of UrbanTwin datasets as both augmentation and standalone training sources. By bridging the sim-to-real gap in roadside lidar perception, UrbanTwin offers a scalable, cost-effective resource for both research and deployment in intelligent transportation systems. The modular simulation framework supports extensibility to new classes, environments, and annotation types. Future work will focus on integrating pedestrian and VRU classes, expanding validation across tracking, segmentation, and sensor fusion tasks, and creating synthetic replicas for additional public lidar datasets as part of the UCF Digital Twin initiative.

### v2

@@ 1 INTRODUCTION
For the V2X-Real case, SECOND is trained solely on the synthetic dataset and evaluated on the real test set of V2X-Real-IC. The results are compared against benchmark models reported in the original V2X-Real paper. The findings show that models trained on synthetic data perform on par with, and in some cases outperform, models trained on real data, setting a new standard for the utility of synthetic datasets in roadside lidar perception tasks.
@@
It is important to note that this work focuses exclusively on the Lidar modality to rigorously evaluate the quality of synthetic point cloud data. While camera data is a valuable component in multi-modal perception systems, it is intentionally excluded from the scope of this study to ensure a controlled analysis of Lidar-centric, sim-to-real transfer.
@@ 2 Related Work
Real-world Roadside Lidar Datasets: Recent research has increasingly focused on using lidar technology for sensing road users from an infrastructure perspective, leading to the development of several roadside lidar datasets. In 2021, the BAAI-VANJEE dataset [7] was introduced, capturing highway and intersection scenarios across various times of day and weather conditions. In 2022, the IPS300+ dataset [4] was released, notable for featuring over 300 road users per frame, primarily due to the high volume of pedestrians and cyclists at a major intersection in Beijing. That same year, the DAIR-V2X dataset [6] became the first large-scale, multi-view dataset designed for vehicle-infrastructure cooperative driving research. Also in 2022, LUMPI [1] was introduced as the largest dataset by frame count, collected over multiple days and weather conditions. Rope3D [8], while primarily focused on monocular 3D object detection using cameras, incorporated lidar data to improve annotation accuracy. In 2023, the TUMTraf dataset [3]—a successor to the A9 dataset [18]—was released, with a focus on non-pedestrian traffic. The V2X-Real dataset [2], introduced in 2024, was the first large-scale dataset to integrate multiple vehicles and smart infrastructure for V2X cooperative perception. This was complemented by other large-scale efforts like RCooper [19], which provides a real-world dataset focused on roadside cooperative perception. Concurrently, advancements in spatio-temporal fusion for multi-agent perception and prediction were explored in works like V2xpnp [20]. That same year, Int2Sec [21] targeted urban intersections using a digital twin perspective, offering 10 scenes annotated via dual roadside lidars. HoloVIC [22] further expanded the space by providing over 100K synchronized frames with 3D object annotations across multiple holographic intersections. Most recently, in 2025, the R-LiViT dataset [5] was introduced, combining lidar, RGB, and thermal imaging from a roadside viewpoint, with a focus on pedestrians and other vulnerable road users.
@@
Synthetic Roadside Lidar Datasets: The recent surge in roadside lidar datasets reflects a growing interest in infrastructure-centric perception. However, creating large-scale, annotated lidar datasets remains an inherently expensive and labor-intensive process, restricting the scalability needed for training safety-critical models [23]. This challenge has accelerated the adoption of simulation, which can speed up research by orders of magnitude [24], and is increasingly motivated by the Digital Twin (DT) paradigm. Foundational work has established how infrastructure-based digital twins can serve as quality-controlled information sources for automated driving functions [25]. As detailed in recent surveys, the DT-ITS framework aims to create virtual replicas of entire transportation ecosystems for full lifecycle management and the validation of perception algorithms [26, 27]. In response to these needs, and to overcome the limitations of real-world data collection, several synthetic lidar datasets have been proposed.
In 2022, SynLidar [28] used a custom simulator to synthesize a large synthetic lidar dataset generated using multiple simulated scenes. In the same year, V2X-Sim and V2X-Set [29] datasets simulated comprehensive multi-agent scenarios to generate synchronized camera and lidar data that included roadside units (RSUs) into the mix for cooperative perception tasks. DOLPHINS [30] dataset included six autonomous driving scenarios involving temporarily synchronized sensor data from an ego vehicle. In 2023, the DeepAccident dataset [31] was released, providing 57K frames that include safety-critical scenarios for evaluating accident prediction models. In 2024, SynthmanticLiDAR [32] introduced a CARLA-based simulator tailored for lidar semantic segmentation, producing annotated point clouds aligned with real-world class distributions. TUMTraf Synthetic (TUMTraf-Syn) Dataset [33] included 24,000 images with depth maps, and 2D and 3D annotations in 10 object categories focusing sim2real monocular 3D object detection. The SCOPE dataset [34], also introduced in 2024, emphasized diversity, featuring realistic sensor models, physically accurate weather conditions, a catalog of over 40 scenarios, up to 24 collaborative agents, and passive traffic.
@@
Sim-to-Real Gap Mitigation Approaches: Synthetic datasets are often created using simulators that model 3D scenes under user-defined conditions. However, the outputs frequently diverge from real data due to unmodeled environmental and sensor-specific effects, resulting in a noticeable sim-to-real domain gap. For Lidar-based 3D object detection, this gap can be particularly severe; recent studies have quantified this effect, showing that state-of-the-art detectors trained exclusively on simulated data can suffer a performance degradation of over 50% in mean Average Precision (mAP) when evaluated on real-world data [35]. Addressing this gap has been a central focus in many recent studies.
@@
3 UrbanTwin Dataset Generation and Description
We present three synthetic roadside lidar datasets: UrbanTwin-LUMPI [40], UrbanTwin-V2X-Real [41], and UrbanTwin-TUMTraf-I [42]. For brevity, these datasets will be referred to using the UT- prefix throughout this article. Each synthetic dataset is designed to replicate the core characteristics of its corresponding real-world counterpart, LUMPI [1], V2X-Real [2], and TUMTraf-I [3].
3.1 Generation Approach
@@ 3.2 Dataset Description
The synthetic datasets contain 6 classes: Car, Van, Bicycle, Motorcycle, Bus, and Truck. Due to the challenging conformation of pedestrian models in CARLA to real data, all pedestrian and VRU classes are omitted for now to maintain high fidelity, and are planned to be added in a future.
@@ 4.1 Structural and Distributional Similarity
Figure 9: Qualitative Comparison of 3D Object Detection Results on Real Datasets Using Models Trained Exclusively on Synthetic Data. Results from two LiDAR-based object detectors, SEED (top) and SECOND (bottom), trained solely on the synthetic UT-LUMPI and UT-V2X-Real datasets, respectively, and evaluated on their corresponding real-world test sets. Each column shows five randomly selected frames from testsets illustrating ground-truth annotations and model predictions. The close alignment between predicted and ground-truth bounding boxes demonstrates strong cross-domain generalization, confirming that models trained on high-fidelity synthetic datasets can accurately detect real-world objects despite the absence of real training data.
@@ 4.2 Utility in Perception Models
To ensure immediate utility and ease of use, all UrbanTwin datasets have been standardized into a consistent format, facilitating practical application in training and testing 3D perception models. The point clouds are unified (per origin and scale) to target datasets, and are provided as .npy files for easy loading. The labels are formatted to be directly compatible with popular OpenPCDet [44] and SemanticKITTI [45] frameworks. Specifically, annotations for 3D object detection (in lbl folder) include object position (x, y, z), size (dx, dy, dz), and heading (yaw angle around the z-axis), with an additional field for the object’s tracking ID. For segmentation tasks, labels in ins_seg and sem_seg folders (for instance and semantic segmentation respectively), follow the standard SemanticKITTI .label format, enabling research in scene understanding.
@@
Case Study - 3D Object Detection: To demonstrate real-world applicability, we evaluate a common perception task, 3D object detection using the synthetic datasets. The objective of this case study is not to benchmark model performance against state-of-the-art methods or to analyze detection architectures in detail. Rather, it serves as a validation exercise to assess the quality of point clouds, labels, and the overall training utility of the UT-datasets.
To this end, we train two off-the-shelf deep 3D object detectors for the car category, SEED [16] and SECOND [17], on the synthetic UT-LUMPI and UT-V2X-Real datasets, respectively, and evaluate their performance on the corresponding real test sets from LUMPI and V2X-Real. To qualitatively assess this transferability, Fig. 9 provides a visual comparison between ground-truth annotations and the predictions from models trained solely on our synthetic data.
@@
V2X-Real vs. UT-V2X-Real: For V2X-Real, we evaluate the generalization capacity of the synthetic dataset by training the SECOND model exclusively on 8,000 synthetic samples from UT-V2X. The trained model is then evaluated on the V2X-Real-IC test set and compared against benchmark results reported in the original V2X-Real paper. This benchmark includes a diverse set of architectures, ranging from single-stage detectors to more advanced attention and transformer-based models like AttFuse and V2X-ViT, providing a robust test of our dataset’s generalization capabilities.
@@ 6 Conclusion & Future Work
This paper introduced three high-fidelity synthetic roadside lidar datasets, UT-LUMPI, UT-V2X-Real, and UT-TUMTraf-I, constructed using a digital-twin-based pipeline to replicate the physical layout and sensor properties of real-world deployments. Unlike generic synthetic datasets, our approach achieves close structural, statistical, and sensory alignment with real data. Through distributional analysis and object detection case studies, we demonstrated that models trained solely on synthetic data generalize well to real-world benchmarks, in some cases outperforming models trained on human-annotated datasets. These results highlight the utility of UrbanTwin datasets as both augmentation and standalone training sources. By bridging the sim-to-real gap in roadside lidar perception, UrbanTwin offers a scalable, cost-effective resource for both research and deployment in intelligent transportation systems. The modular simulation framework supports extensibility to new classes, environments, and annotation types. Future work will focus on integrating pedestrian and other Vulnerable Road User (VRU) classes, whose inclusion is critical for safety-related ITS research [47]. To overcome current simulation limitations, we plan to integrate realistic human models using technologies like Unreal Engine’s MetaHuman framework [48] and source naturalistic motion from large-scale motion capture libraries (e.g. AMASS [49]). We will also explore cross-site generalization, expanding validation across tracking, segmentation, and sensor fusion tasks, and continue creating synthetic replicas for additional public lidar datasets as part of the UCF Digital Twin initiative.